OpenStudy (anonymous):

Trig/ Pre Cal/identies Am I on the right path? Please do not give answer but guide me. Thank you. The problem \[ \frac{1}{\sin x + 1} + \frac{1}{\csc x +1} \] \[ \frac{\csc x + 1 + \sin x + 1}{( \sin x + 1)( \csc x + 1)}\] Am I on the right path?

OpenStudy (anonymous):

it should be \[ \frac{1}{\sin x + 1} + \frac{1}{\csc x +1} = 1 \] \[ \frac{\csc x + 1 + \sin x + 1}{( \sin x + 1)( \csc x + 1)} = 1\]

Nnesha (nnesha):

well first write csc in terms of sin :-) that will be easy :-)

Nnesha (nnesha):

\[ \frac{1}{\sin x + 1} + \frac{1}{\rm \color{reD}{\frac{1}{sinx} +1}} = 1 \] like this solve red part first

OpenStudy (anonymous):

So \[ \frac{\frac{1}{sin x} x + 1 + \sin x + 1}{( \sin x + 1)( \frac{1}{sin x} x + 1)}\]

Nnesha (nnesha):

better to gget rid of fraction first :-)

OpenStudy (anonymous):

so do I times the numerator and denominator by sin??

Nnesha (nnesha):

\[ \frac{1}{\sin x + 1} + \frac{1}{\rm \color{reD}{\frac{1}{sinx} +1}} = 1 \] \[\large\rm \color{reD}{\frac{ 1 }{ \sin x } +1}\] multiply by sinx/1 do you know how to find common denominator ? just for the red part

OpenStudy (anonymous):

For the red part the common denominator is sin x and 1. For the red we have \[ 1+ \sin x \] corect?

Nnesha (nnesha):

for the red part  common denominator is sin x